A freely falling body travels how much %in its 5th set? 
 

Grade:11

1 Answers

Arun
5 years ago

As you need to find the distance covered in fifth second, you should find the distance covered in six seconds and five seconds and subtract them.

As per the equation s=ut+1/2(at^2)

u=0, a=g (acceleration due to gravity) and t = time

Distance covered in 6 seconds (let’s call it s6)= 1/2 (36g)

Distance covered in 5 seconds (let’s call it s5)= 1/2 (25g)

therefore answer is s6-s5= 11g/2 (put value of g as 10 or 9.8 as per you accuracy requirements and solve to get answer).